a piggy bank contains $5.85 in nickels and quarters. if there are 65 coins in all, how many of each type coin is there?
please help me

Answer by stanbon(57361)

a piggy bank contains $5.85 in nickels and quarters. if there are 65 coins in all, how many of each type coin is there?
----------------------------------------------
Quantity Equation: n + q = 65
Value Equation::: 5n + 25q = 585 cents
--------------------------------------------
Multiply thru the 1st Eq. by 5 to set up the system for elimination:
5n + 5q = 5*65
5n + 25q = 585
--------------------
Subract the 1st from the 2nd to solve for "q":
20q = 260
q = 13 (# of quarters)
-----------------
Substitute in n + q = 65 to solve for "n"
n + 13 = 65
n = 52 (# of nickels)
========================================
Cheers,
Stan H.